The table below describes results from random selection of eight live births from eight different sets of parents. The random variable x represents the number of girls among the eight live births. Is 6 a significantly high number of girls out of 8 live births?


1
Expert's answer
2022-03-09T10:13:46-0500

"P(X\ge 6)=P(X=6)+P(X=7)+P(X=8)="

"=C^6_80.5^60.5^2+C^7_80.5^70.5^1+C^8_80.5^80.5^0=0.1445."

Since this probability is greater than 0.05, 6 is not a significantly high number of girls out of 8 live births.